The accompanying summary data on total cholesterol level (mmol/l) was obtained from a sample of Asian postmenopausal women who were vegans and another sample of such women who were omnivores.

Diet Sample Size Sample Mean Sample SD
Vegan 89 5.30 1.05
Omnivore 92 5.75 1.10

Calculate a 99% CI for the difference between population mean total cholesterol level for vegans and population mean total cholesterol level for omnivores. (Use μVeganμOmnivore. Round your answers to three decimal places.)

  ,

(mmol/l)

Interpret the interval.

We are 99% confident that the true average cholesterol level for vegans is less than that of omnivores by an amount within the confidence interval.We are 99% confident that the true average cholesterol level for vegans is greater than that of omnivores by an amount within the confidence interval.    We cannot draw a conclusion from the given information.We are 99% confident that the true average cholesterol level for vegans is greater than that of omnivores by an amount outside the confidence interval.
Answer #1

The standard error here is computed as:

SE = \sqrt{\frac{s_1^2}{n_1} + \frac{s_2^2}{n_2}} = \sqrt{\frac{1.05^2}{89} + \frac{1.1^2}{92}} = 0.1598

For n1 + n2 - 2 = 179 degrees of freedom, we have from t distribution tables here:
P( t179 < 2.604) = 0.995

Therefore, P( -2.604 < t179 < 2.604) = 0.99

Therefore the confidence interval here is obtained as:

\bar X_1 - \bar X_2 \pm t^*SE

5.3 - 5.75 \pm 2.604*0.1598

- 0.45 \pm 0.4161

(- 0.8661, -0.0339)

This is the required 99% confidence interval for the difference in population means here.

b) As the whole confidence interval here lies below 0, therefore we are 99% confident here that the true average cholesterol level for vegans is greater than that of omnivores by an amount within the confidence interval